Here is the problem: The length of a rectangle is 4 meters longer than the width. If the area is 8 square meters, find the rectangle's dimensions. Round to the nearest tenth of a meter. I keep arriving at 4 m. by 1.5 m and that is NOT correct...Thank you in advance for your help! Amber
Answer by mananth(16949)

Here is the problem: The length of a rectangle is 4 meters longer than the width. If the area is 8 square meters, find the rectangle's dimensions. Round to the nearest tenth of a meter. I keep arriving at 4 m. by 1.5 m and that is NOT correct...Thank you in advance for your help! Amber
let the width be x meters
length will be x+4 meters
x(x+4)= 8
x^2+4x-8=0
x1= -4+sqrt(16+32) / 2
x1= -4 +sqrt(48) /2
= 1.46 = 1.5
x2= -5.46
so 1.5 is the width
X+4 is the length = 1.5 +4 = 5.5 meters
5.5*1.46 = 8.03 square ,meters
Ananth